# Mitigating Content Effects on Reasoning in Language Models Through Fine-Grained Activation Steering

### 06 Mar 2026 v2

@@ 1 Introduction
Large language models (LLMs) possess advanced natural and common-sense reasoning capabilities but are prone to content effects, i.e., systematic biases where prior knowledge and believability of content influence logical inference (Bertolazzi et al. 2024a; Lampinen et al. 2024). For example, an LLM may incorrectly judge a logically invalid syllogism as valid if its content aligns with common-sense knowledge (e.g., “All students read; some readers are professors; therefore some students are professors”), mirroring human content biases (Lampinen et al. 2024). Such behaviour violates the requirement of formal reasoning, where validity should depend only on logical form, not content. Recent studies have documented these content-based reasoning failures (Bertolazzi et al. 2024a), showing that models find factually believable premises easier to “prove” and struggle with abstract or counter-intuitive ones (Lampinen et al. 2024). This undermines LLMs’ reliability on formal reasoning, particularly logic-oriented tasks highlighted by Bertolazzi et al. (2024b).
On the other hand, prompting strategies alone are insufficient to eliminate content effects. Chain-of-thought (CoT) prompting and related methods (Wei et al. 2022; Kojima et al. 2022) can improve reasoning. However, biases often persist in the generated explanations, and models may still arrive at content-biased conclusions even when “thinking aloud” (Ranaldi et al. 2025). In fact, (Bertolazzi et al. 2024b) finds that while CoT and fine-tuning boost accuracy on logical deductions, they do not fully remove biases like content believability effects. Similarly, neuro-symbolic approaches have been proposed to improve robustness in formal reasoning with LLMs (Quan et al. 2024, 2025b, 2025a; Pan et al. 2023; Lyu et al. 2023). However, they introduce the complexity of integrating LLMs with external symbolic solvers.
@@
Expanding on previous work (Bertolazzi et al. 2024a; Kim et al. 2025; Wysocka et al. 2025), we generate a synthetic dataset leveraging known syllogistic arguments, considering the intersection of plausible/implausible and formally valid/invalid arguments. The dataset includes over 16k arguments generated by instantiating 24 abstract syllogistic schemes with the support of Wordnet (Miller 1995).
@@ 2 Background
Recent research has demonstrated that LLMs exhibit content effects in formal reasoning tasks, mirroring cognitive biases that may align or differ from those observed in humans (Dasgupta et al. 2022; Kim et al. 2025; Mondorf and Plank 2024; Seals and Shalin 2024; Wysocka et al. 2025; Eisape et al. 2024). These effects arise when the semantic plausibility of a prompt influences the model’s reasoning process, often leading to correct conclusions for plausible statements and systematic errors for implausible but logically valid ones (Bertolazzi et al. 2024a).
(Dasgupta et al. 2022) first showed that LLMs perform better on reasoning tasks when the content of the problem aligns with world knowledge. In their experiments, models were significantly more accurate on syllogistic tasks when the conclusions were semantically plausible, even when this plausibility conflicted with the actual logical validity of the argument. This indicates a bias toward material reasoning (reasoning grounded in semantic associations) rather than formal reasoning (reasoning based strictly on logic).
Further work by (Bertolazzi et al. 2024a) systematically evaluated LLMs on a broad suite of syllogisms and found that performance dropped sharply for arguments that contradicted commonsense knowledge. This reliance on content plausibility suggests that LLMs are susceptible to semantic interference, failing to uphold the norms of formal logic when they conflict with prior knowledge. (Seals et al. 2023) and (Agarwal et al. 2024) also emphasized this discrepancy, showing that even the most capable frontier models tend to conflate logical validity with plausibility. To the best of our knowledge, this is the first work investigating how content effects can be reduced through activation steering techniques (Subramani et al. 2022; Hernandez et al. 2024; Zou et al. 2023; Turner et al. 2023; Li et al. 2023; Zhao et al. 2024).
@@ 3 Methodology
Our goal is to investigate and mitigate content effects in LLMs, i.e., systematic biases where semantic plausibility influences logical reasoning. To this end, we design a controlled syllogistic reasoning task, leveraging 24 abstract syllogistic schemes automatically instantiated through taxonomic knowledge from external knowledge bases (Sec. 3.1) and apply activation-level steering techniques to modulate model behavior toward formal validity assessment (Sec. 3.2). Further, we identify the limitations of current state-of-the-art steering methods, and propose a more fine-grained steering approach (Sec. 3.2). Figure 1 provides a high-level overview of the methodology.
@@ 3.1 Formal Syllogistic Reasoning
Inspired by recent work (Lample and Charton 2019; Bertolazzi et al. 2024a; Wysocka et al. 2025; Kim et al. 2025), we evaluate formal reasoning in LLMs through syllogistic arguments. We formalise the task of syllogistic reasoning as a binary classification problem, where the objective is to determine the formal validity of a syllogism, whether its conclusion follows logically from its premises, irrespective of the plausibility of the content. Specifically, the model is expected to predict VALID or INVALID based solely on the logical form in the syllogism structure $\mathcal{S}$.
@@
We construct a dataset of approximately 16,000 syllogistic arguments in English to systematically analyze content effects. Each argument instantiates one of the 24 formal syllogistic schemas (details in the supplementary material) and is explicitly varied along dimensions of formal validity and semantic plausibility.
The data generation process begins with the formalization of syllogistic structures in first-order logic (FOL), following prior work on logical reasoning datasets (Bertolazzi et al. 2024a; Wysocka et al. 2025). Each logical schema is then converted into natural language templates such as: All A are B, All B are C, All A are C.
@@ Appendix B Syllogistic Schemes
Here is the list of syllogistic schemes adopted to construct the dataset. The schemes are aligned with the ones used in previous work on syllogistic reasoning with LLMs (Bertolazzi et al. 2024a).

### 01 Apr 2026 v3

@@ 1 Introduction
Large language models (LLMs) possess advanced natural and common-sense reasoning capabilities but are prone to content effects – i.e., systematic biases where prior knowledge and believability of content influence logical inference (Bertolazzi et al. 2024; Lampinen et al. 2024). For example, an LLM may incorrectly judge a logically invalid syllogism as valid if its content aligns with common-sense knowledge (e.g., “All students read; some readers are professors; therefore some students are professors”), mirroring human content biases (Lampinen et al. 2024). Such behaviour violates the requirement of formal reasoning, where validity should depend only on logical form, not content. Recent studies have documented these content-based reasoning failures (Bertolazzi et al. 2024), showing that models find factually believable premises easier to “prove” and struggle with abstract or counter-intuitive ones (Lampinen et al. 2024). This undermines LLMs’ reliability on formal reasoning, particularly logic-oriented tasks highlighted by Bertolazzi et al. (2024).
On the other hand, prompting strategies alone are insufficient to eliminate content effects. Chain-of-thought (CoT) prompting and related methods (Wei et al. 2022; Kojima et al. 2022) can improve reasoning. However, biases often persist in the generated explanations, and models may still arrive at content-biased conclusions even when “thinking aloud” (Ranaldi et al. 2025). In fact, (Bertolazzi et al. 2024) finds that while CoT and fine-tuning boost accuracy on logical deductions, they do not fully remove biases like content believability effects. Similarly, neuro-symbolic approaches have been proposed to improve robustness in formal reasoning with LLMs (Quan et al. 2024, 2025b, 2025a; Pan et al. 2023; Lyu et al. 2023). However, they introduce the complexity of integrating LLMs with external symbolic solvers.
@@
Expanding on previous work (Bertolazzi et al. 2024; Kim et al. 2025; Wysocka et al. 2025), we generate a synthetic dataset leveraging known syllogistic arguments, considering the intersection of plausible/implausible and formally valid/invalid arguments. The dataset includes over 16k arguments generated by instantiating abstract syllogistic schemes with the support of Wordnet (Miller 1995).
@@ 2 Background
Recent research has demonstrated that LLMs exhibit content effects in formal reasoning tasks, mirroring cognitive biases that may align or differ from those observed in humans (Lampinen et al. 2024; Kim et al. 2025; Mondorf and Plank 2024; Seals and Shalin 2024; Wysocka et al. 2025; Eisape et al. 2024). These effects arise when the semantic plausibility of a prompt influences the model’s reasoning process, often leading to correct conclusions for plausible statements and systematic errors for implausible but logically valid ones (Bertolazzi et al. 2024).
(Lampinen et al. 2024) first showed that LLMs perform better on reasoning tasks when the content of the problem aligns with world knowledge. In their experiments, models were significantly more accurate on syllogistic tasks when the conclusions were semantically plausible, even when this plausibility conflicted with the actual logical validity of the argument. This indicates a bias toward material reasoning (reasoning grounded in semantic associations) rather than formal reasoning (reasoning based strictly on logic).
Further work by (Bertolazzi et al. 2024) systematically evaluated LLMs on a broad suite of syllogisms and found that performance dropped sharply for arguments that contradicted commonsense knowledge. This reliance on content plausibility suggests that LLMs are susceptible to semantic interference, failing to uphold the norms of formal logic when they conflict with prior knowledge. (Seals and Shalin 2024) also emphasized this discrepancy, showing biases in LLMs’ deductive reasoning. To the best of our knowledge, this is the first work investigating how content effects can be reduced through activation steering techniques (Subramani et al. 2022; Hernandez et al. 2024; Zou et al. 2023; Turner et al. 2023; Li et al. 2023; Zhao et al. 2024).
@@ 3 Methodology
Our goal is to investigate and mitigate content effects in LLMs, i.e., systematic biases where semantic plausibility influences logical reasoning. To this end, we design a controlled syllogistic reasoning task, leveraging abstract syllogistic schemes automatically instantiated through taxonomic knowledge from external knowledge bases (Sec. 3.1) and apply activation-level steering techniques to modulate model behavior toward formal validity assessment (Sec. 3.2). Further, we identify the limitations of current state-of-the-art steering methods, and propose a more fine-grained steering approach (Sec. 3.2). Figure 1 provides a high-level overview of the methodology.
@@ 3.1 Formal Syllogistic Reasoning
Inspired by recent work (Bertolazzi et al. 2024; Wysocka et al. 2025; Kim et al. 2025; Lampinen et al. 2024), we evaluate formal reasoning in LLMs through syllogistic arguments. We formalise the task of syllogistic reasoning as a binary classification problem, where the objective is to determine the formal validity of a syllogism, whether its conclusion follows logically from its premises, irrespective of the plausibility of the content. Specifically, the model is expected to predict VALID or INVALID based solely on the logical form in the syllogism structure $\mathcal{S}$.
@@
We construct a dataset of approximately 16,000 syllogistic arguments in English to systematically analyze content effects. Each argument instantiates one of the formal syllogistic schemas (details in the supplementary material) and is explicitly varied along dimensions of formal validity and semantic plausibility.
The data generation process begins with the formalization of syllogistic structures in first-order logic (FOL), following prior work on logical reasoning datasets (Bertolazzi et al. 2024; Wysocka et al. 2025). Each logical schema is then converted into natural language templates such as: All A are B, All B are C, All A are C.
@@ Appendix B Syllogistic Schemes
Here is the list of syllogistic schemes adopted to construct the dataset. The schemes are aligned with the ones used in previous work on syllogistic reasoning with LLMs (Bertolazzi et al. 2024).